OpenStudy (anonymous):

-2/5x - 9 < 9/10 show steps please

OpenStudy (anonymous):

ummmm what lol

OpenStudy (anonymous):

this inequality has no solution.

OpenStudy (anonymous):

can you show me how you got that tho?

OpenStudy (anonymous):

@binks

OpenStudy (anonymous):

To set the left-hand side of the inequality equal to 0, move all the expressions to the left-hand side. -2/5x-9-9/10<0 To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 10x. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. -2/5x . 2/2.-9.10x/10x-9/10.x/x<0 Complete the multiplication to produce a denominator of 10x in each expression -4-90x-9x/10x <0 Combine all like terms in the numerator -99x+4/10x <0 Find all the values where the expression switches from negative to positive by setting each factor equal to 0 and solving. 10x=0 (99x+4)=0 Solve for each factor to find the values where the absolute value expression goes from negative to positive x=0 x=-4/99 Since there are no numbers that fall within the interval, this inequality has no solution

OpenStudy (anonymous):

im guessing the . means multiplication

OpenStudy (anonymous):

oh yeah sorry

OpenStudy (anonymous):

haha its ok I just guessed that is what it meant